HUBBLE SPACE TELESCOPE DAILY REPORT #1341 PERIOD COVERED: 0000Z (UTC) 03/06/95 - 0000Z (UTC) 03/07/95 Daily Status Report as of 066/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/48 6144 (FOC "F/48 High Voltage Anomaly Investigation) The Faint Object Camera (f/48) was used to make a dark and internal flat calibration images. This is part of a test to try to determine the cause of the high-voltage discharge in the f/48 intensifier. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 5472 (The Collimation and Propagation of Jets from Young Stars) The WF/PC-2 was used to make observations of the variable star HL-Tau to study its environment. The intent is to study the outflow (jet) from this star and determine the nature of the mechanism that produces the observed structures. The observations were executed as planned, and no problems were reported. 1.4 Completed Two Sets of GHRS/1 5276 (GHRS Side 1 Detector Calibration : Cycle 4) The High Resolution Spectrograph (Side One) was used to conduct a standard detector test that will be used as part of the routine calibration for the HRS. The observations were completed as planned, and no anomalies were reported. 1.5 Completed WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C430.0 as part of a continuum survey and to study the galaxy's morphology. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed FOS/RD 5692 (A Study of the Arcsecond Gravitational Lens 1422+231: Cycle 4 Carryover.) The Faint Object Spectrograph (Red Side) was used to make observations of the four components of gravitationally lensed system QSO1422P23. The observations were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 6142 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Smaller Set Part II) The WF/PC-2 was used to make Earth Flat calibration images for the ultraviolet and narrow-band filters to correct for the telescope's illumination pattern. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were completed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 1070 (High Resolution Imaging of Radio Galaxies at Large Redshift) The WF/PC-2 was used to make deep images of radio galaxy 3C356 to study the evolution of its stellar population and to look for star formation. The observations were completed as planned, and no anomalies were reported. 1.11 Completed Two Sets of GHRS/1 5595 (GHRS ECH-A Wavelength Monitor: Cycle 4) The High Resolution Spectrograph was used to make wavelength spectral lamp calibration observations to measure the instrumental characteristics of the Echelle-A spectrograph. The observations were completed as planned, and no problems were reported. 1.12 Completed FOS/RD 6141 (The Aftermath of SL-9 and Jupiter) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of the planet Jupiter to study the aftermath of the impact of the comet SL-9. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed GHRS/1 5413 (The Local ISM Flow Within the Heliosphere) The High Resolution Spectrograph was used to make observations to study the flow of neutral atoms in the interstellar medium by observing diffuse sky background in Lyman Alpha. A wavelength calibration image was as taken. The spectral calibration lamp was left on at 066/0615Z, at the end of observation #20 (HSTAR #5148). The lamp was turned off by ground command. No other anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Significant HSTARs: None 2.4 Operations Notes: There are no new operational problems to report. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-